OpenStudy (anonymous):

3/7 divided by 3 = 1/4 divided by 5 = 1/8 divided by 6 =

OpenStudy (zepp):

divide by anything mean multiplying by its inverse, so (3/7)/3 could be changed to (3/7)*(1/3) There you have some fractions ;)

OpenStudy (anonymous):

3/7 * 1/3 = 1/7 1/4 * 1/5 = 1/20 1/8 * 1/6 = 1/48

OpenStudy (anonymous):

yeah wat he said